OpenStudy (anonymous):

A petroleum company has two different sources of crude oil. The first source provides crude oil that is 30% hydrocarbons, and the second one provides crude oil that is 70% hydrocarbons. In order to obtain 180 gallons of crude oil that is 60% hydrocarbons, how many gallons of crude oil must be used from each of the two sources? first source: _ gallons second source: _ gallons

OpenStudy (anonymous):

another copy and paste fail

OpenStudy (anonymous):

The first source provides crude oil that is ????? hydrocarbons,

OpenStudy (anonymous):

I fixed it.

OpenStudy (anonymous):

put \(x\) as the amount of 30% crude, then since the total is 180 the amount of 70% crude is \(180-x\) \[.3x+.7(180-x)=.6\times 180\] solve for \(x\)

OpenStudy (anonymous):

probably easier to get rid of annoying decimals and solve \[3x+7(180-x)=6\times 180\]

OpenStudy (anonymous):

thank you(:

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

give me a min to think lol